OpenStudy (anonymous):

Mary is experimenting with chemicals in the laboratory. She mixes a solution that is 10% acid with a solution that is 25% acid. How much of the 10% acid solution will be needed to make 50 liters of solution that is 20% acid?

OpenStudy (cwrw238):

the amount of pure acid in the 20% solution = 0.2 * 50 = 10 liters let the volume of 10% solution be x liters then amount of pure acid in this is 0.1x liters and amount of acid in 25% solution will be (50 - x)*0.25 liters so equating these volumes of pure acid: 0.1x + 0.25(50 - x) = 10 0.1x + 12.5 - 0.25x = 10 2.5 = 0.15x x = 2.5 / 0.15 = 16.33 liters of 10% solution

OpenStudy (cwrw238):

to be more accurate 16.33.... liters ( 16 1/3)

OpenStudy (anonymous):

thank you cwrw238

OpenStudy (cwrw238):

yw